# 1]DAMO Academy, Alibaba Group

### v1

@@ 1 Introduction
In recent years, Multi-modal Large Language Models (MLLMs) [62, 77] have experienced rapid development, leading to the emergence of models such as Gemini [58] and GPT-4o [46] that can handle image and even video inputs. These MLLMs are attracting increasing attention from researchers due to their powerful contextual understanding [15] and generalization [82] abilities. Researchers in embodied intelligence are also beginning to explore the use of MLLMs as the brains of robots [22, 26], enabling them to perceive the real world through visual inputs like humans do. However, the current mainstream MLLMs are trained on extensive internet images and lack the foundational visual cognition to match the physical world [13, 74].
Some works have begun exploring how MLLMs can be applied to ego-centric embodied scenarios. Models like Exo2Ego [79] and EgoLM [23] enhance the understanding of ego-centric dynamic environment interactions. SpatialVLM [7] and SpatialRGPT [10] focus on addressing spatial understanding challenges within embodied contexts. However, these approaches are challenging to directly implement in physical robots to perform complex tasks. The main limitations are as follows:
@@
    Absence of video-based coherent spatial awareness: For humans, spatial cognition arises from continuous visual perception [48]. Current methods in spatial intelligence [80, 64] primarily focus on single or discrete images, lacking the capacity for spatial understanding in high-continuity videos. For example, in Task 1 Step 4 in Fig. 1, the absolute distance between the teddy bear and the pillow requires a spatial scale concept derived from the entire video to be properly inferred.
Thus, we propose RynnEC, an embodied cognitive MLLM designed to enhance robotic understanding of the physical world. As illustrated in Fig. 1, RynnEC is a large video understanding model whose visual encoder and foundational parameters are derived from VideoLLaMA3 [76]. To enable flexible visual interaction, we incorporate an encoder and decoder specifically for region masks in videos, allowing RynnEC to achieve precise instance-level comprehension and grounding.
@@
Regrettably, the development of embodied cognition models has been slow due to a lack of ego-centric videos and high-quality annotations. Efforts such as Multi-SpatialMLLM [64], Spatial-MLLM [61], and SpaceR [47] leverage open-source datasets with comprehensive 3D point cloud and annotations to generate training data. However, in an era of scarce 3D annotations [24, 38], this approach cannot achieve rapid and cost-effective expansion of data scale. Hence, we propose a data generation pipeline that transforms ego-centric RGB videos into embodied cognition question-answering datasets. This pipeline begins with instance segmentation from videos and diverges into two branches: one generating object cognition data and the other producing spatial cognition data. Ultimately, data from both branches are integrated into a comprehensive embodied cognition dataset. From over 200 households, we collect more than 20,000 egocentric videos. A subset from ten households is manually verified and balanced to create RynnEC-Bench, a fine-grained embodied cognition benchmark encompassing 22 tasks in object and spatial cognition.
Extensive experiments demonstrate that RynnEC significantly outperforms both general [46, 4, 87] and task-specific [72, 75, 57] MLLMs in cognitive abilities within embodied scenarios, showcasing scalable application potential. Additionally, we observe notable advantages in multi-task training with RynnEC and identify preliminary signs of emergence in more challenging embodied cognition tasks. Finally, we highlight the potential of RynnEC in facilitating robots to undertake large-scale, long-range tasks.
@@ 2.1 MLLMs for Video Understanding
Early MLLMs primarily relied on sparse sampling and simple connectors, such as MLPs [33, 2, 39] and Q-Formers [78, 28], to integrate visual representation with large language models. Subsequently, to tackle the problem of long video understanding, [81] directly expanded the context window of language models, while [83] introduced pooling in the spatial and temporal dimensions to compress the number of video tokens. As the need for more fine-grained understanding emerged, some studies (VideoRefer [75], DAM [32] and PAM [35]) employed region-level feature encoders enabling video MLLMs to accept masked inputs and comprehend the semantic features of objects within the masks. Although these video MLLMs have demonstrated superior capabilities in high-level semantic capture and temporal modeling, they lack robust physical-world comprehension in egocentric embodied scenarios.
@@ 2.2 Embodied Scene Understanding Benchmarks
Some studies [54, 29, 21] have begun to explore leveraging MLLMs to assist robots in solving embodied tasks. However, determining whether these MLLMs possess the ability to understand and interact with the physical world is challenging. Consequently, several benchmarks have emerged to evaluate the capability of MLLMs to perceive the physical world. OpenEQA [41] and IndustryEQA [30] focus on several key competencies in home and industrial settings, respectively, and manually designed open-vocabulary questions. VSI-Bench [69] centers on assessing the spatial cognitive abilities of MLLMs. STI-Bench [31] introduces more complex kinematic (e.g. velocity) problems. ECBench [13] systematically categorizes embodied cognitive abilities into static environments, dynamic environments, and overcoming hallucinations, offering a comprehensive evaluation across 30 sub-competencies. While these benchmarks encompass a wide range of abilities, they are unable to assess more fine-grained, region-level understanding capabilities in embodied scenarios. Compared to purely textual question-answering, region-level visual interaction can more accurately refer to targets in the complex real world.
@@ 2.3 Improving MLLMs for Embodied Cognition
The aforementioned embodied benchmarks have highlighted the cognitive limitations of current MLLMs in embodied scenarios. Consequently, some studies have started to investigate diverse strategies for enhancing MLLMs’ understanding of the physical world. GPT4Scene [49] improves MLLMs’ consistent global scene understanding by explicitly adding instance marks between video frames. SAT [53] explores multi-frame dynamic spatial reasoning in simulated environments. Spatial-MLLM [61], Multi-SpatialMLLM [64], and SpaceR [47] leverage 3D datasets with detailed annotations (e.g., ScanNet [70]) to construct the suite of spatial-intelligence tasks introduced in VSI-Bench. In contrast, our data generation pipeline based on RGB videos yields more realistic and scalable training data. More importantly, RynnEC is designed not just to handle selected capabilities in embodied scenarios, but to cover a broad swath of the world cognition required for embodied task execution under a single paradigm.
@@ 3.1.1 Video Collection and Instance Segmentation
Previous works [37, 60] adopted a strategy of designing separate data generation processes for each task type, leading to limited data reusability and continuity. We aim to create a lineage among different types of foundational data to reduce unnecessary redundancy in data generation. Therefore, this paper proposes a mask-centric embodied cognition QA generation pipeline. This pipeline initiates with the generation of object masks from video instance segmentation within a scene. First, Qwen2.5-VL [4] observes the raw video and outputs an object list containing the names of all entity categories in the scene. Utilizing this object list, Grounding DINO 1.5 [55] detects objects in key frames at one-second intervals. SAM2 [51] assists in segmenting and tracking the objects detected by Grounding DINO 1.5 during the intervening one-second interval. To ensure consistency of instance IDs, the tracking results of old instances are compared with the segmentation results of newly detected instances at key frames. If an instance is found to have overlapping masks (IOU > 0.5), it retains the ID of the old tracking instance. Due to the performance limitations of Grounding DINO 1.5, newly detected object instances may already have appeared in preceding frames yet were missed. Thus, SAM2 conducts a reverse four-second instance tracking for each new object in key frames, thereby achieving full lifecycle instance tracking. In total, we obtain 1.14 million video instance masks from all the egocentric videos.
@@ 3.1.2 Object QA Generation
Due to the limitation of SAM2 in consistent object tracking in egocentric videos, the same instance may be assigned multiple IDs if the instance appears intermittently in the video. We employ an object category filtering method that limits each video to a maximum of two instances per object category, thereby minimizing duplicate instances. The presence of multiple video segments per house leads to repeated occurrences of certain salient objects, causing a pronounced long-tail distribution. We downsample object categories that occur frequently to prevent extreme object distribution. After the aforementioned filtering, the cue image sets of retained instances are input into Qwen2.5-VL [4], generating object caption and object comprehension QA through various prompts. It is noteworthy that in the object comprehension QA, counting QA task is particularly unique and requires specially designed prompts. Subsequently, based on each instance’s caption and QAs, Qwen3 [67] generates two types of referring expressions: simple referring expressions and situational referring expressions. Simple referring expressions identify objects through a combination of features such as spatial location and category. Situational referring expressions establish a task scenario, requiring the model to infer the instance needed by the user within this context. Each type of QA undergoes manual filtering post-output to ensure data quality. Detailed prompts are provided in the Appendix A.2.
@@ 3.1.3 Spatial QA Generation
Unlike object QA, spatial QA requires more precise 3D information concerning the global scene context. Therefore, we utilize MASt3R-SLAM [44] to reconstruct 3D point clouds from RGB videos and obtain camera extrinsic parameters. Subsequently, by projecting 2D pixel points to 3D coordinates, the segmentation of each instance in the video can be mapped onto the point cloud. However, it is important to note that the world coordinate system established by MASt3R-SLAM for the 3D point cloud is not aligned with the floor. Therefore, the Random Sample Consensus (RANSAC) [19] algorithm is implemented to identify inlier points for plane fitting through ten iterative executions. In each iteration, the detected planar surface and its inliers are removed from the point cloud for subsequent plane detection. Given that the initial camera pose was approximately horizontal but not perpendicular to the ground, the ground plane is selected based on minimal angular deviation between its normal vector and the initial camera Y-axis orientation. The point cloud is then aligned to ensure orthogonality between the world coordinate Z-axis and the detected ground plane.
@@
Building on insights from prior works [61, 47], we recognize that spatial cognition tasks are highly challenging. Therefore, in addition to constructing a large-scale video-based spatial QA dataset, we also develop a relatively simpler image-based spatial QA dataset. This combination of tasks with varying levels of difficulty is intended to improve learning efficiency and enhance model robustness. Specifically, we collect 500k indoor images from 39k houses. Leveraging the single-image-to-3D reconstruction and calibration methods from SpatialRGPT [10], we obtain the 3D spatial relationships between objects in each image. We then select tasks from the video-based spatial cognition set that can also be addressed via single images, and design corresponding QA templates. The format of the image-based spatial QA is kept consistent with that of the video-based spatial QA.
@@ 3.2.3 Evaluation Framework
The questions are categorized into three types based on the nature of their answers: numerical questions, textual questions, and segmentation questions. For numerical questions such as distance estimation and direction estimation, we directly use the formula to calculate the precise indicators. For scale-related questions, Mean Relative Accuracy (MRA) [69, 16] is used to calculate the scores. Specifically, given a model’s prediction $\hat{y}$, ground truth $y$, and a confidence threshold $\theta$, relative accuracy is calculated by considering $\hat{y}$ correct if the relative error rate, defined as $|\hat{y}-y|/y$, is less than $1-\theta$. As single-confidence-threshold accuracy only considers relative error within a narrow scope, the MRA averages the relative accuracy across a range of confidence thresholds $\mathcal{C}=\{0.5,\,0.55,\,\ldots,\,0.95\}$:
@@
For segmentation evaluation, prior work [72, 65] typically reports the $\mathcal{J}\&\mathcal{F}$ measure, combining region-overlap ($\mathcal{J}$) and boundary-accuracy ($\mathcal{F}$) scores. However, the conventional frame-averaged $\mathcal{J}\&\mathcal{F}$ treats empty frames (i.e., frames with no ground-truth mask) in a binary manner: if any predicted mask appears, the frame score is set to 0; otherwise it is set to 1. This evaluation method fails to account for the actual size of erroneous masks in empty frames, which can have a significant impact on embodied segmentation tasks. To address this, we propose the Global IoU metric, defined as
@@ 3.3 RynnEC Architecture
Foundational Vision-Language Model. We ultilize VideoLLaMA3-Image [76] as the foundational vision-language model for RynnEC, which consists of three main modules: a Vision Encoder, the Projector and the Large Language Model (LLM). For the vision encoder, we use VL3-SigLIP-NaViT [76], which leverages an any-resolution vision tokenization strategy to flexibly encode images of varying resolutions. As the LLM, we employ Qwen2.5-1.5B-Instruct [66] and Qwen2.5-7B-Instruct [66], enabling scalable trade-offs between performance and computational cost.
Region Encoder. Egocentric videos often feature cluttered scenes with similar objects that are difficult to distinguish using linguistic cues alone. To address this, we introduce a dedicated object encoder for specific object representation. This facilitates more precise cross-modal alignment during training and enables intuitive, fine-grained user interaction at inference time. Following  [73, 75], we use a simple yet efficient MaskPooling for object tokenization, followed by a two-layer projector to align object features with LLM embedding space. During training, object masks spanning multiple frames in a video are utilized to achieve accurate representations. At inference, the encoder offers flexibility, operating effectively with either single-frame or multi-frame object masks.
Mask Decoder. Accurate object localization is critical for egocentric video understanding. To incorporate robust visual grounding capabilities without degrading the model’s pretrained performance, we fine-tune the LLM with LoRA. Our mask decoder is based on the architecture of SAM2 [51], which has demonstrated strong generalization capabilities and prior knowledge in purely visual segmentation tasks. For a given video and the instruction, we adpot a [SEG] token as a specifical token to trigger mask generation for the corresponding visual region. To facilitate this process, an additional linear layer is introduced to align the [SEG] token with SAM2’s feature space.

### v2

@@ 1 Introduction
In recent years, Multi-modal Large Language Models (MLLMs) Wu et al. [2023], Zhang et al. [2024a] have experienced rapid development, leading to the emergence of models such as Gemini Team et al. [2024] and GPT-4o OpenAI et al. [2024] that can handle image and even video inputs. These MLLMs are attracting increasing attention from researchers due to their powerful contextual understanding Doveh et al. [2025] and generalization Zhang et al. [2024c] abilities. Researchers in embodied intelligence are also beginning to explore the use of MLLMs as the brains of robots Han et al. [2025b], Jin et al. [2024], enabling them to perceive the real world through visual inputs like humans do. However, the current mainstream MLLMs are trained on extensive internet images and lack the foundational visual cognition to match the physical world Dang et al. [2025], Yuan et al. [2025b].
Some works have begun exploring how MLLMs can be applied to ego-centric embodied scenarios. Models like Exo2Ego Zhang et al. [2025b] and EgoLM Hong et al. [2025] enhance the understanding of ego-centric dynamic environment interactions. SpatialVLM Chen et al. [2024a] and SpatialRGPT Cheng et al. [2024a] focus on addressing spatial understanding challenges within embodied contexts. However, these approaches are challenging to directly implement in physical robots to perform complex tasks. The main limitations are as follows:
@@
    Absence of video-based coherent spatial awareness: For humans, spatial cognition arises from continuous visual perception Pasqualotto and Proulx [2012]. Current methods in spatial intelligence Zhang et al. [2025c], Xu et al. [2025] primarily focus on single or discrete images, lacking the capacity for spatial understanding in high-continuity videos. For example, in Task 1 Step 4 in Fig. 1, the absolute distance between the teddy bear and the pillow requires a spatial scale concept derived from the entire video to be properly inferred.
Thus, we propose RynnEC, an embodied cognitive MLLM designed to enhance robotic understanding of the physical world. As illustrated in Fig. 1, RynnEC is a large video understanding model whose visual encoder and foundational parameters are derived from VideoLLaMA3 Zhang et al. [2025a]. To enable flexible visual interaction, we incorporate an encoder and decoder specifically for region masks in videos, allowing RynnEC to achieve precise instance-level comprehension and grounding.
@@
Regrettably, the development of embodied cognition models has been slow due to a lack of ego-centric videos and high-quality annotations. Efforts such as Multi-SpatialMLLM Xu et al. [2025], Spatial-MLLM Wu et al. [2025a], and SpaceR Ouyang et al. [2025] leverage open-source datasets with comprehensive 3D point cloud and annotations to generate training data. However, in an era of scarce 3D annotations Hou et al. [2025], Lyu et al. [2024], this approach cannot achieve rapid and cost-effective expansion of data scale. Hence, we propose a data generation pipeline that transforms ego-centric RGB videos into embodied cognition question-answering datasets. This pipeline begins with instance segmentation from videos and diverges into two branches: one generating object cognition data and the other producing spatial cognition data. Ultimately, data from both branches are integrated into a comprehensive embodied cognition dataset. From over 200 households, we collect more than 20,000 egocentric videos. A subset from ten households is manually verified and balanced to create RynnEC-Bench, a fine-grained embodied cognition benchmark encompassing 22 tasks in object and spatial cognition.
Extensive experiments demonstrate that RynnEC significantly outperforms both general OpenAI et al. [2024], Bai et al. [2025], Zhu et al. [2025] and task-specific Yuan et al. [2025a, c], Team et al. [2025] MLLMs in cognitive abilities within embodied scenarios, showcasing scalable application potential. Additionally, we observe notable advantages in multi-task training with RynnEC and identify preliminary signs of emergence in more challenging embodied cognition tasks. Finally, we highlight the potential of RynnEC in facilitating robots to undertake large-scale, long-range tasks.
@@ 2.1 MLLMs for Video Understanding
Early MLLMs primarily relied on sparse sampling and simple connectors, such as MLPs Lin et al. [2023], Ataallah et al. [2024], Maaz et al. [2023] and Q-Formers Zhang et al. [2023], Li et al. [2024b], to integrate visual representation with large language models. Subsequently, to tackle the problem of long video understanding, Zhang et al. [2024b] directly expanded the context window of language models, while Zhang et al. [2024d] introduced pooling in the spatial and temporal dimensions to compress the number of video tokens. As the need for more fine-grained understanding emerged, some studies (VideoRefer Yuan et al. [2025c], DAM Lian et al. [2025] and PAM Lin et al. [2025]) employed region-level feature encoders enabling video MLLMs to accept masked inputs and comprehend the semantic features of objects within the masks. Although these video MLLMs have demonstrated superior capabilities in high-level semantic capture and temporal modeling, they lack robust physical-world comprehension in egocentric embodied scenarios.
@@ 2.2 Embodied Scene Understanding Benchmarks
Some studies Ren et al. [2024a], Li et al. [2024c], Han et al. [2025a] have begun to explore leveraging MLLMs to assist robots in solving embodied tasks. However, determining whether these MLLMs possess the ability to understand and interact with the physical world is challenging. Consequently, several benchmarks have emerged to evaluate the capability of MLLMs to perceive the physical world. OpenEQA Majumdar et al. [2024] and IndustryEQA Li et al. [2025a] focus on several key competencies in home and industrial settings, respectively, and manually designed open-vocabulary questions. VSI-Bench Yang et al. [2025c] centers on assessing the spatial cognitive abilities of MLLMs. STI-Bench Li et al. [2025b] introduces more complex kinematic (e.g. velocity) problems. ECBench Dang et al. [2025] systematically categorizes embodied cognitive abilities into static environments, dynamic environments, and overcoming hallucinations, offering a comprehensive evaluation across 30 sub-competencies. While these benchmarks encompass a wide range of abilities, they are unable to assess more fine-grained, region-level understanding capabilities in embodied scenarios. Compared to purely textual question-answering, region-level visual interaction can more accurately refer to targets in the complex real world.
@@ 2.3 Improving MLLMs for Embodied Cognition
The aforementioned embodied benchmarks have highlighted the cognitive limitations of current MLLMs in embodied scenarios. Consequently, some studies have started to investigate diverse strategies for enhancing MLLMs’ understanding of the physical world. GPT4Scene Qi et al. [2025] improves MLLMs’ consistent global scene understanding by explicitly adding instance marks between video frames. SAT Ray et al. [2024] explores multi-frame dynamic spatial reasoning in simulated environments. Spatial-MLLM Wu et al. [2025a], Multi-SpatialMLLM Xu et al. [2025], and SpaceR Ouyang et al. [2025] leverage 3D datasets with detailed annotations (e.g., ScanNet Yeshwanth et al. [2023]) to construct the suite of spatial-intelligence tasks introduced in VSI-Bench. In contrast, our data generation pipeline based on RGB videos yields more realistic and scalable training data. More importantly, RynnEC is designed not just to handle selected capabilities in embodied scenarios, but to cover a broad swath of the world cognition required for embodied task execution under a single paradigm.
@@ 3.1.1 Video Collection and Instance Segmentation
Previous works Luo et al. [2025], Wang et al. [2024] adopted a strategy of designing separate data generation processes for each task type, leading to limited data reusability and continuity. We aim to create a lineage among different types of foundational data to reduce unnecessary redundancy in data generation. Therefore, this paper proposes a mask-centric embodied cognition QA generation pipeline. This pipeline initiates with the generation of object masks from video instance segmentation within a scene. First, Qwen2.5-VL Bai et al. [2025] observes the raw video and outputs an object list containing the names of all entity categories in the scene. Utilizing this object list, Grounding DINO 1.5 Ren et al. [2024b] detects objects in key frames at one-second intervals. SAM2 Ravi et al. [2024] assists in segmenting and tracking the objects detected by Grounding DINO 1.5 during the intervening one-second interval. To ensure consistency of instance IDs, the tracking results of old instances are compared with the segmentation results of newly detected instances at key frames. If an instance is found to have overlapping masks (IOU > 0.5), it retains the ID of the old tracking instance. Due to the performance limitations of Grounding DINO 1.5, newly detected object instances may already have appeared in preceding frames yet were missed. Thus, SAM2 conducts a reverse four-second instance tracking for each new object in key frames, thereby achieving full lifecycle instance tracking. In total, we obtain 1.14 million video instance masks from all the egocentric videos.
@@ 3.1.2 Object QA Generation
Due to the limitation of SAM2 in consistent object tracking in egocentric videos, the same instance may be assigned multiple IDs if the instance appears intermittently in the video. We employ an object category filtering method that limits each video to a maximum of two instances per object category, thereby minimizing duplicate instances. The presence of multiple video segments per house leads to repeated occurrences of certain salient objects, causing a pronounced long-tail distribution. We downsample object categories that occur frequently to prevent extreme object distribution. After the aforementioned filtering, the cue image sets of retained instances are input into Qwen2.5-VL Bai et al. [2025], generating object caption and object comprehension QA through various prompts. It is noteworthy that in the object comprehension QA, counting QA task is particularly unique and requires specially designed prompts. Subsequently, based on each instance’s caption and QAs, Qwen3 Yang et al. [2025a] generates two types of referring expressions: simple referring expressions and situational referring expressions. Simple referring expressions identify objects through a combination of features such as spatial location and category. Situational referring expressions establish a task scenario, requiring the model to infer the instance needed by the user within this context. Each type of QA undergoes manual filtering post-output to ensure data quality. Detailed prompts are provided in the Appendix A.2.
@@ 3.1.3 Spatial QA Generation
Unlike object QA, spatial QA requires more precise 3D information concerning the global scene context. Therefore, we utilize MASt3R-SLAM Murai et al. [2025] to reconstruct 3D point clouds from RGB videos and obtain camera extrinsic parameters. Subsequently, by projecting 2D pixel points to 3D coordinates, the segmentation of each instance in the video can be mapped onto the point cloud. However, it is important to note that the world coordinate system established by MASt3R-SLAM for the 3D point cloud is not aligned with the floor. Therefore, the Random Sample Consensus (RANSAC) Fischler and Bolles [1981] algorithm is implemented to identify inlier points for plane fitting through ten iterative executions. In each iteration, the detected planar surface and its inliers are removed from the point cloud for subsequent plane detection. Given that the initial camera pose was approximately horizontal but not perpendicular to the ground, the ground plane is selected based on minimal angular deviation between its normal vector and the initial camera Y-axis orientation. The point cloud is then aligned to ensure orthogonality between the world coordinate Z-axis and the detected ground plane.
@@
Building on insights from prior works Wu et al. [2025a], Ouyang et al. [2025], we recognize that spatial cognition tasks are highly challenging. Therefore, in addition to constructing a large-scale video-based spatial QA dataset, we also develop a relatively simpler image-based spatial QA dataset. This combination of tasks with varying levels of difficulty is intended to improve learning efficiency and enhance model robustness. Specifically, we collect 500k indoor images from 39k houses. Leveraging the single-image-to-3D reconstruction and calibration methods from SpatialRGPT Cheng et al. [2024a], we obtain the 3D spatial relationships between objects in each image. We then select tasks from the video-based spatial cognition set that can also be addressed via single images, and design corresponding QA templates. The format of the image-based spatial QA is kept consistent with that of the video-based spatial QA.
@@ 3.2.3 Evaluation Framework
The questions are categorized into three types based on the nature of their answers: numerical questions, textual questions, and segmentation questions. For numerical questions such as distance estimation and direction estimation, we directly use the formula to calculate the precise indicators. For scale-related questions, Mean Relative Accuracy (MRA) Yang et al. [2025c], Everingham et al. [2010] is used to calculate the scores. Specifically, given a model’s prediction $\hat{y}$, ground truth $y$, and a confidence threshold $\theta$, relative accuracy is calculated by considering $\hat{y}$ correct if the relative error rate, defined as $|\hat{y}-y|/y$, is less than $1-\theta$. As single-confidence-threshold accuracy only considers relative error within a narrow scope, the MRA averages the relative accuracy across a range of confidence thresholds $\mathcal{C}=\{0.5,\,0.55,\,\ldots,\,0.95\}$:
@@
For segmentation evaluation, prior work Yuan et al. [2025a], Yan et al. [2024] typically reports the $\mathcal{J}\&\mathcal{F}$ measure, combining region-overlap ($\mathcal{J}$) and boundary-accuracy ($\mathcal{F}$) scores. However, the conventional frame-averaged $\mathcal{J}\&\mathcal{F}$ treats empty frames (i.e., frames with no ground-truth mask) in a binary manner: if any predicted mask appears, the frame score is set to 0; otherwise it is set to 1. This evaluation method fails to account for the actual size of erroneous masks in empty frames, which can have a significant impact on embodied segmentation tasks. To address this, we propose the Global IoU metric, defined as
@@ 3.3 RynnEC Architecture
Foundational Vision-Language Model. We ultilize VideoLLaMA3-Image Zhang et al. [2025a] as the foundational vision-language model for RynnEC, which consists of three main modules: a Vision Encoder, the Projector and the Large Language Model (LLM). For the vision encoder, we use VL3-SigLIP-NaViT Zhang et al. [2025a], which leverages an any-resolution vision tokenization strategy to flexibly encode images of varying resolutions. As the LLM, we employ Qwen2.5-1.5B-Instruct Yang et al. [2024] and Qwen2.5-7B-Instruct Yang et al. [2024], enabling scalable trade-offs between performance and computational cost.
Region Encoder. Egocentric videos often feature cluttered scenes with similar objects that are difficult to distinguish using linguistic cues alone. To address this, we introduce a dedicated object encoder for specific object representation. This facilitates more precise cross-modal alignment during training and enables intuitive, fine-grained user interaction at inference time. Following Yuan et al. [2024, 2025c], we use a simple yet efficient MaskPooling for object tokenization, followed by a two-layer projector to align object features with LLM embedding space. During training, object masks spanning multiple frames in a video are utilized to achieve accurate representations. At inference, the encoder offers flexibility, operating effectively with either single-frame or multi-frame object masks.
Mask Decoder. Accurate object localization is critical for egocentric video understanding. To incorporate robust visual grounding capabilities without degrading the model’s pretrained performance, we fine-tune the LLM with LoRA. Our mask decoder is based on the architecture of SAM2 Ravi et al. [2024], which has demonstrated strong generalization capabilities and prior knowledge in purely visual segmentation tasks. For a given video and the instruction, we adpot a [SEG] token as a specifical token to trigger mask generation for the corresponding visual region. To facilitate this process, an additional linear layer is introduced to align the [SEG] token with SAM2’s feature space.